Gas in a piston-cylinder apparatus is heated with 150.0 J while simultaneously decreasing its volume from 1.500 L to 0.500 L at a constant pressure of 1.000 bar. Calculate the total energy change in Joules for this process.

Question 6 options:

0 (the work and heat contributions cancel out)

+50.0 J

-150.0 J

-50.0 J

+250.0 J
Answer #1

use:

W = -Pext(Vf-Vi)

Given:

P= 1.0 bar

= (1.0/1.01325) atm

= 0.9869 atm

Vi = 1.5 L

Vf = 0.5 L

put values in above expression

W = -Pext(Vf-Vi)

W = -0.9869 atm (0.5 L - 1.5 L)

W = 0.9869 atm.L

W = 0.9869 * 101.33 J

W = 100.0 J

Given:

Q = 150.0 J

Use:

Delat E = Q+W

= 150.0 J = 100.0 J

= 250.0 J

Answer: 250.0 J